OpenStudy (anonymous):

Can you guys help me please T.T ... Given that the quadratic equation x^2-2x-5=0 has 2 different roots, and a second quadratic equation has 2 roots, each of which is 2 less than the corresponding root of the given quadratic equation. If the second quadratic equation is x^2+ax+b=0, find the value of a and b.

OpenStudy (anonymous):

wow. i think you actually have to find the roots. we can do that fairly easily

OpenStudy (anonymous):

root are 2 and -5

OpenStudy (anonymous):

\[x^2-2x=5\] \[(x-1)^2=5+1=6\] \[x--=\pm\sqrt{6}\] \[x=1\pm\sqrt{6}\]

OpenStudy (anonymous):

so the two roots are \[1-\sqrt{6}\] and \[1-\sqrt{6}\] subtract 2 from each get \[-1-\sqrt{6}\] and \[-1+\sqrt{6}\]

OpenStudy (anonymous):

now multiply out and see what you get in the equation \[(x-(-1-\sqrt{6}))(x-(-1+\sqrt{6})\] which looks hard but in fact is very easy

OpenStudy (anonymous):

*expression

OpenStudy (anonymous):

first \[x^2\] last \[(-1-\sqrt{6})(-1+\sqrt{6})=1-6=-5\] hmm just like the orignal

OpenStudy (anonymous):

"outer, inner" \[(1-\sqrt{6})x+(1+\sqrt{6})x=2x\]

OpenStudy (anonymous):

so equation is \[x^2+2x-5=0\]

OpenStudy (anonymous):

they want you to find the a and b value

OpenStudy (radar):

You have them,\[ax ^{2}+2x-5=0\] a=1 b=2